OpenStudy (anonymous):

Solve by the substituion method. 2x - 7 = 12 and 8x + 60 = y. What is the solution of the equation as an ordered pair?

OpenStudy (saifoo.khan):

an soo... bunny is back!

OpenStudy (anonymous):

FINALLY!

OpenStudy (anonymous):

get first equation in terms of x

OpenStudy (saifoo.khan):

check your 1st eq.

OpenStudy (anonymous):

Subtract 2x from both sides?

OpenStudy (saifoo.khan):

in your 1st eq there is no y.

OpenStudy (anonymous):

oh crap...that is supposed to be 7y

OpenStudy (saifoo.khan):

-7y?

OpenStudy (anonymous):

supposed to be 2x - 7y = 12

OpenStudy (anonymous):

K. still get it in terms of x... or y...

OpenStudy (saifoo.khan):

try to find the ordered pair. i got it. ;)

OpenStudy (saifoo.khan):

gand, she knws how ti solve it.

OpenStudy (anonymous):

already? showoff!

OpenStudy (saifoo.khan):

lol, srsly.

OpenStudy (saifoo.khan):

want the answer?

OpenStudy (anonymous):

YES

OpenStudy (saifoo.khan):

but no, i wont giv it.. try to solev it yourself.

OpenStudy (saifoo.khan):

solve*

OpenStudy (anonymous):

Sorry dumbunny, thought u wanted to know how

OpenStudy (anonymous):

I do

OpenStudy (anonymous):

okay so for the first equation, I should subtract 8x from both sides?

OpenStudy (anonymous):

no, I meant 2x from both sides

OpenStudy (anonymous):

which variable u want 2 isolate?

OpenStudy (anonymous):

I think y

OpenStudy (saifoo.khan):

why y?

OpenStudy (saifoo.khan):

JK.

OpenStudy (anonymous):

K. then -2x is a very good start :)

OpenStudy (anonymous):

so then would I have 7y = 12-2x

OpenStudy (anonymous):

7y = 10x

OpenStudy (anonymous):

-7y

OpenStudy (saifoo.khan):

why dont use elimination method>

OpenStudy (anonymous):

and you can't combine the 12 and the 2x

OpenStudy (anonymous):

so I am back to 7y = 12-2x

OpenStudy (anonymous):

-7y=12x+2

OpenStudy (anonymous):

What are ya'll doing? lol

OpenStudy (anonymous):

I can't use the elimination method because it asks me to use the substitution method.

OpenStudy (anonymous):

Isolating y when y is already isolated? o_O

OpenStudy (anonymous):

divide each side by -7

OpenStudy (saifoo.khan):

Sj, spamming.

OpenStudy (saifoo.khan):

Oh, okay.. proceed.. @bunny.

OpenStudy (saifoo.khan):

LOL @sj

OpenStudy (anonymous):

okay so I have 7y = 12 - 2x and I need to divide both sides by -7?

OpenStudy (anonymous):

u have -7y=12-2x

OpenStudy (anonymous):

and u divide each side by -7

OpenStudy (anonymous):

OHHHH I see... Didn't see where she wrote it was supposed to be 7y ^_^

OpenStudy (anonymous):

so I would then have y = 12-2x/-7...I think I am confused

OpenStudy (anonymous):

y=12-2x ----- -7

OpenStudy (saifoo.khan):

best option, help her on whiteboard.

OpenStudy (anonymous):

really sorry bunny, but I got to get off. I think the others can help u

OpenStudy (anonymous):

okay, thanks

OpenStudy (saifoo.khan):

come here Bunny: http://www.twiddla.com/588961

OpenStudy (anonymous):

dumbunny a better way to look at it that's easier (at least for me) is [y={12\over-7} -{2\over-7}x\]

OpenStudy (anonymous):

Except if the equation actually worked...

OpenStudy (anonymous):

\[y={12\over-7}-{2\over-7}x\]